OpenStudy (anonymous):

Factor the following binomial. y^2-25 I got (y-5)(y+5) is this right

OpenStudy (accessdenied):

Using difference of squares: a^2 - b^2 = (a - b)(a + b) y^2 - 25 = y^2 - 5^2 = (y - 5)(y + 5) is correct. If you wanted to check, you could multiply back through by FOIL. :)

OpenStudy (anonymous):

in my book the answer is (y+5)(y-5)

OpenStudy (anonymous):

Multiplication is commutative...meaning a * b = b * a...so your answer and the book's answer are the same.

OpenStudy (anonymous):

oh ok thanks @mtbender74